40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Sep 13 14:12:03 2022, Yehonathan, Update, BHD, Trying LO phase locking again Screenshot_2022-09-13_17-41-33.png
    Reply  Thu Sep 15 21:12:53 2022, Paco, Update, BHD, LO phase "dc" control lophase_oltf.pdflophase_noise_spectra.pdf
       Reply  Tue Sep 27 10:50:11 2022, Paco, Update, BHD, calibrated LO phase noise lo_phase_asd.pdf
          Reply  Wed Sep 28 16:37:26 2022, Paco, Update, BHD, calibrated LO phase noise; update lo_phase_asd.pdf
             Reply  Wed Sep 28 21:54:08 2022, Paco, Update, BHD, calibrated LO phase noise; update dcpd_phase_asd.pdflo_phase_asd.pdf
                Reply  Fri Sep 30 20:18:55 2022, Paco, Update, BHD, LO phase noise with different actuation points BHDC_asd_actuation_point.png
                   Reply  Mon Oct 3 15:19:05 2022, Paco, Update, BHD, LO phase noise and control after violin mode filters filters_c1sus_lo1_lsc.pngBHDC_asd_act.pngboosted_lo_phase_control.png
Message ID: 17171     Entry time: Mon Oct 3 15:19:05 2022     In reply to: 17167
Author: Paco 
Type: Update 
Category: BHD 
Subject: LO phase noise and control after violin mode filters 

[Anchal, Paco]

We started the day by taking a spectrum of C1:HPC-LO_PHASE_IN1, the BHD error point, and confirming the absence of 268 Hz peaks believed to be violin modes on LO1. We then locked the LO phase by actuating on LO2, and AS1. We couldn't get a stable loop with AS4 this morning. In all of these trials, we looked to see if the noise increased at 268 Hz or its harmonics but luckily it didn't. We then decided to add the necessary output filters to avoid exciting these violin modes. The added filters are in the C1:SUS-LO1_LSC bank, slots FM1-3 and comprise bandstop filters at first, second and third harmonics observed previously (268, 536, and 1072 Hz); bode plots for the foton transfer functions are shown in Attachment #1. We made sure we weren't adding too much phase lag near the UGF (~ 1 degree @ 30 Hz).

We repeated the LO phase noise measurement by actuating on LO1, LO2 and AS1, and observe no noise peaks related to 268 Hz this time. The calibrated spectra are in Attachment #2. Now the spectra look very similar to one another, which is nice. The rms is still better when actuating with AS1.


[Paco]

After the above work ended, I tried enabling FM1-3 on the C1:HPC_LO_PHASE control filters. These filters boost the gain to suppress noise at low frequencies. I carefully enabled them when actuating on LO1, and managed to suppress the noise by another factor of 20 below the UGF of ~ 30 Hz. Attachment #3 shows the screenshot of the uncalibrated noise spectra for (1) unsupressed (black, dashed), (2) suppressed with FM4-5 (blue, solid), and (3) boosted FM1-5 suppression (red).


Next steps:

  • Compare LO-ITMY and LO-ITMX single bounce noise spectra and MICH.
  • Compare DC locking scheme versus BH55 once it's working.
Attachment 1: filters_c1sus_lo1_lsc.png  28 kB  Uploaded Mon Oct 3 17:06:33 2022  | Hide | Hide all
filters_c1sus_lo1_lsc.png
Attachment 2: BHDC_asd_act.png  239 kB  Uploaded Mon Oct 3 17:08:33 2022  | Hide | Hide all
BHDC_asd_act.png
Attachment 3: boosted_lo_phase_control.png  36 kB  Uploaded Mon Oct 3 17:25:16 2022  | Hide | Hide all
boosted_lo_phase_control.png
ELOG V3.1.3-